OpenStudy (anonymous):

Which equation would best help solve the following problem: Mark launches a toy rocket with an initial vertical velocity of 25 feet per second from a platform that is at ground level. How long will the toy rocket be in the air?

OpenStudy (anonymous):

1 second.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

We may misunderstand the question. It might be this one: Until the rocket lands, how long has it been in the air? If so, solve this question this way. The equation of the object's trajectory: y = -g*t^2/2 + 25t = t(-gt/2 + 25). The rocket lands at t = 0 (at base) and a t seconds given by: g*t/2 = 25 -> t = 50/9.81 = 5.09 sec